OpenStudy (anonymous):

Choose the Standard Form of the equation y = - x - 4.

OpenStudy (shamim):

x+y+4=0

OpenStudy (shamim):

m i right

OpenStudy (anonymous):

the equation is y=-1/3x-4

OpenStudy (texaschic101):

standard form is Ax + By = C your equation is : y = -1/3x - 4 (multiply the equation by 3 to get rid of the fraction) 3y = -1x - 12 x + 3y = -12 is your equation in standard form

OpenStudy (texaschic101):

any questions ?

OpenStudy (anonymous):

This question has two parts. Which of the following is not a step used when graphing the equation y = 1/2-3 using the slope-intercept method? Explain the correct order of steps needed to graph this equation. Plot the y-intercept (0, -3). Draw a line through the two points. Count up 1 and right 2 from the y-intercept. Plot the x-intercept (6, 0).

OpenStudy (anonymous):

Ontan lass VIejas :))

OpenStudy (texaschic101):

I am sorry...I am no good at graphing